OpenStudy (anonymous):

Write an equation of an ellipse centered at the origin with the height 8 units and width 16 units.

OpenStudy (jdoe0001):

well... hm one sec

OpenStudy (jdoe0001):

is centered at the origin... so hmmm  so... now recall that \(\bf \cfrac{(x-{\color{brown}{ h}})^2}{{\color{purple}{ a}}^2}+\cfrac{(y-{\color{blue}{ k}})^2}{{\color{purple}{ b}}^2}=1 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad vertices\ ({\color{brown}{ h}}\pm a, {\color{blue}{ k}})\) so.. .what do you think is our "h" and "k" and our "a" and "b" components?

TheSmartOne (thesmartone):

\(\sf\Large \frac{x^2}{\left( \frac{width}{2} \right)^2} + \frac{y^2}{\left(\frac{height}{2}\right)^2}=1\)

TheSmartOne (thesmartone):

If the height was greater than the width, then (height/2)^2 would be under x^2 and (width/2)^2 would be under y^2

TheSmartOne (thesmartone):

@l_o_w_key It would be nice if you could tell us where you are stuck at, or if you have understood what you have been told. :)

OpenStudy (anonymous):

(0,16)(8,0)?

OpenStudy (anonymous):

forget that. one sec

OpenStudy (jdoe0001):

ahemm ... and so :)

TheSmartOne (thesmartone):

:)

OpenStudy (jdoe0001):

just to clarify on the graph above, it's wider, and the width runs over the x-axis thus the "a" component goes below the numerator with the "x" variable

OpenStudy (anonymous):